base of a rectangle=10 m and the height is 7m.What is the perimeter dilated to half the size?
I appreciate your help.

Found 2 solutions by TakeATuition.com, nycsharkman:
Answer by TakeATuition.com(57)

SOLUTION -
Base=10 m
height=7 m

now for ba new rectangle
base=5m
height=3.5m
perimeter=2(l+b)
=2(5+3.5)
=2*8.5
=17 m
FOR MORE HELP VISIT:www.TakeATuition.com

Answer by nycsharkman(136)

A dilation is a transformation that produces an image that is the same shape as the original, but is a different size.
A dilation used to create an image larger than the original is called an enlargement. A dilation used to create an image smaller than the original is called a reduction.
The scale factor here is 1/2.
Perimeter = 2L + 2W
Length = base
Width = height
P = 2(10m) + 2(7m)
P = 20m + 14m
P = 34m
To find the dilated perimeter, we multiply 34m by the scale factor 1/2.
P = 34m times 1/2
P = 17m
I hope this helps.